OpenStudy (anonymous):

Find the indicated limit, if it exists. HOW do I do this?! Lim x-->5 f(x),f(x){5-x x<5 {8 x=5 {x+3 x>5 0 8 3 The limit does not exist

OpenStudy (anonymous):

@amistre64 ?

OpenStudy (amistre64):

it looks like you are given a piecewise function

OpenStudy (amistre64):

the limit will only exist of 5-x = x+3, when x=5 otherwise the limit for the left and right of 5 do not match up

OpenStudy (amistre64):

be sure that the question is not asking for a one-sided limit like:\[x\to 5^+\]

OpenStudy (amistre64):

if we were to graph these it would look something like this: 

OpenStudy (amistre64):

notice: if we come into x=5 from the left we approach 0 if we come into x=5 from the right we approach 8 since 8 does not equal 0, the limit of the function does not exist ....

OpenStudy (anonymous):

I am plugging in 5 for X in each one and adding/subtracting the total to get my answer, is that right? and what do I do for the 8 x=5?

OpenStudy (amistre64):

the limit of a function has no bearing on the value of the function at the limit point .... the limit of a function only cares if the value from the left and the right are equal

OpenStudy (amistre64):

we see that the value at x=5 is 8 ... the limit doesnt care what this value is ... the value from the left as x approaches 5 is defined by (x-5) the value from the right as x approaches 5 is defined by (x+3) x-5 has no bad calculations in it so itll just be: 5-5 = 0 x+3 has no bad calculations in it so itll just be: 5+3 = 8 since 8 is not equal to 0, the limit does not exist

OpenStudy (anonymous):

So i'm not plugging in 5 for x?

OpenStudy (amistre64):

your not caring that the function defines f(5) = 8. otherwise, yes. you are comparing the definition of the function from left and right as the values of x approach 5.

OpenStudy (amistre64):

 if the values are not the same, then there is no limit